# MARS: Meaning-Aware Response Scoring for Uncertainty Estimation in Generative LLMs

### v2

@@ 1 Introduction
Uncertainty Estimation (UE) is a well-studied problem in classification scenarios, especially in the computer vision domain Lakshminarayanan et al. (2017); Gal and Ghahramani (2016); Shen et al. (2021). The proposed UE methods in classification tasks, which rely on the class probabilities, are not directly applicable to generative LLMs due to the auto-regressive generative structure of LLMs Malinin and Gales (2021), which implies that LLMs generate text sequentially by predicting each subsequent word based on the combined context of all preceding words. This process differs significantly from classification tasks, where the output is typically a single label or a set of labels assigned to an entire input, without the sequential and context-accumulating nature of generative LLMs. Recent work Malinin and Gales (2021), formalizes how to adapt popular UE methods developed for classification tasks to the context of generative LLMs. They propose using length-normalized scoring to estimate the likelihood of a sequence generated by the model, and the subsequent works Kuhn et al. (2023); Lin et al. (2023); Chen and Mueller (2023) utilize that idea of length-normalized scoring.
@@ 2 Background
In this section, we will go over probability-based UE methods that our work built on, for a detailed discussion on related works, refer to Appendix A.
@@ 2.3 Length-Normalized Scoring
One of the key issues with using sequence probability $P(\mathbf{s}|\mathbf{x},\theta)$ as a proxy for $P(y|x,\theta)$ lies in its tendency to decrease as the sequence length increases. To overcome this issue, Malinin and Gales (2021) uses a length-normalized scoring function instead of sequence probability. [^1] A scoring function $K$ takes two inputs: the predicted probability $p$ of an event and its actual outcome $o$, and returns a numerical score Gneiting and Raftery (2007). Length-normalized scoring $\tilde{P}(\mathbf{s}|\mathbf{x},\theta)$ is defined as follows:
@@ 2.4 Entropy-Based UE for Generative LLMs
where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters. Similar to length-normalized scoring in (5), semantic entropy lacks theoretical justification, yet shows empirical success Kuhn et al. (2023). In Appendix B, we provide a theoretical explanation for these heuristic design choices.
@@ 3.3 Importance Function Design
Our initial approach for assigning importance coefficients to tokens assumes their semantic independence even though tokens often exhibit semantic inter-dependencies. For example, in the sentence “Hamlet is written by William Shakespeare,” tokens “William” and “Shakespeare” are intrinsically linked. Treating such tokens independently ignores linguistic nuances, so we refine our methodology. Instead of masking tokens individually, we mask tokens at the phrase level (details in Appendix C.1). This approach acknowledges and preserves the inherent semantic relationships between closely related tokens, resulting in a more accurate and contextually aware assessment of token importance.
In particular, a response $\textbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$ is composed of phrases $\{h_{1},h_{2},\ldots,h_{K}\}$, where each token $s_{l}$ belongs to a phrase $h_{k}$. We mask phrases one by one and find the importance coefficient of each phrase with BEM model. To translate phrase-level importance coefficients into token-level coefficients, we distribute the importance score to all tokens in the phrase equally. We summarize the enhanced algorithm in Appendix C.2. Further, in Section 4.3, we show that allocating importance score only to the most uncertain token within a phrase also yields comparable results.
@@
The necessity of performing a separate neural network pass for each phrase to determine its importance score increases the computational load of the proposed approach. Additionally, detecting phrases themselves requires another neural network pass, further increasing the computational complexity. To address these challenges, we have developed a BERT-like neural network model with 110M parameters (a significantly smaller model compared to LLMs). This model is capable of performing both tasks simultaneously for a given sequence in a single neural network pass: it identifies phrases within the generated text and their importance scores (see Figure 3). This dual-functionality substantially reduces the computational cost, making the algorithm more efficient and scalable. For detailed model architecture and performance metrics, please refer to Appendix C.
@@ 4.1 Experimental Design
We use three closed-book Question-Answer (QA) datasets for evaluation: TriviaQA Joshi et al. (2017), Natural Questions Kwiatkowski et al. (2019), and WebQA Chang et al. (2022). We give further details in Appendix D.
@@
Following previous works Malinin and Gales (2021); Kuhn et al. (2023), we use Area Under the Receiver Operating Characteristic Curve (AUROC) score for our UE performance metric. AUROC quantifies a method’s ability to distinguish between two classes by plotting the true positive rate against the false positive rate for various threshold values. AUROC score is the area under this curve, ranging from 0 to 1. Higher AUROC score indicates a superior performance, while a score of 0.5 implies a random chance. In our case, ground truth is the correctness [^2] We use GPT-3.5-turbo for evaluating the correctness of the model, as in Lin et al. (2023); Chen and Mueller (2023). of the model response to the question and the prediction is the output of an UE method.
@@ 4.2 Main Results
We present our detailed results in Table 1. Figure 4 illustrates the average AUROC scores for each method and model across three distinct datasets. Upon closer examination of the results, it becomes apparent that the application of MARS consistently improves all baseline methods across various datasets and models. Specifically, MARS yields improvements of up to 5.8 points for Confidence, 6.24 points for Entropy, and 1.51 points for SE.
@@
Figure 4: Average AUROC scores of UE methods over 3 different datasets for various LLMs. The improvement of MARS on top of baselines is shown in green.
@@ 4.5 UE in Medical QA Dataset
Next, we evaluate the UE methods using a medical QA dataset. Publicly available medical QA datasets typically fall into two categories: those with multiple-choice questions Pal et al. (2022); Kotonya and Toni (2020); Jin et al. (2021) and those without clear ground truths Zhu et al. (2019, 2020). To tackle this, we create a subset from the MedMCQA multiple-choice dataset Pal et al. (2022), selecting questions that can be answered objectively without multiple choices. For this, we collaborate with medical professionals to ensure the accuracy and relevance of the selected questions, yielding a dataset of 415 samples. We use AdaptLLM’s Medicine-Chat Cheng et al. (2023), a medical-domain adapted LLaMA-2-Chat-7B model [^3] https://huggingface.co/AdaptLLM/medicine-chat. To evaluate the correctness of model-generated responses, we leverage GPT-4 OpenAI (2023) and assess response validity in the medical domain.
@@ 8 Acknowledgement
This work is supported in part by a research gift from USC-Amazon Center on Secure and Trusted Machine Learning [^4] https://trustedai.usc.edu . Also, we are thankful to MD. Mehmet Sahin Ulukanli and MD. Burak Yaldiz for their invaluable collaboration and expertise in the development of the Medical QA dataset.
@@ Appendix A Related Works
Appendix B Conceptualizing the Response Semantics in Generative LLM Probabilities
In classification tasks, the class probability reflects the model’s confidence in assigning a specific class to an input. It is inherently tied to the semantics of the class. For instance, if a well-calibrated classifier gives a 75% probability to the label “cat” for a given question, it suggests a 75% likelihood that the answer of the question is indeed a cat. This output probability is not only a numerical value; it conveys a semantic understanding of the image content as a cat. However, previously proposed length-normalized scoring and semantic entropy definitions for generative LLMs (Sections 2.3 and 2.4) do not directly correspond to the semantics of the LLM generation. Moreover, they are not proper probability and entropy definitions, lacking theoretical background. Hence, we propose a new random variable that is directly related to the semantics of the output and provide a justification for the heuristic decisions of the previous works Kuhn et al. (2023); Malinin and Gales (2021).
Let $Y$ be a random variable with arbitrary dimension corresponding to the meaning of the sequences generated by an LLM parametrized with $\theta$. The values of $Y$ can be the set of all possible meanings of generated sequences in a given context. Formally, the set is $\{g(\textbf{s},\textbf{x})\}_{\textbf{s}\in\mathcal{S},\textbf{x}\in\mathcal{X}}$, where $g(\cdot)$ is the meaning function that takes generated sentence s and context x as inputs and returns the meaning as output. By defining the properties of the meaning function $g(\cdot)$ and the distribution of $Y$, we can rationalize the heuristic design choices made by previous works.
Malinin and Gales (2021) considers $g(\cdot)$ as a one-to-one function which means that each unique sentence in the given context corresponds to different meanings. In this case, the distribution of $Y$ is defined by using the length-normalized scoring of the generated sequences. More formally
Figure 6: In classification tasks, output probabilities give the probability of the semantic meaning. In the case of generative LLMs, probabilities of semantic meaning are unknown. Thus, we propose an alternative probability distribution MARS for generative LLMs.
On the other hand, Kuhn et al. (2023) claims different sequences can have equal meaning. By considering $g(\cdot)$ as a many-to-one function, we can write their proposal with the new meaning random variable $Y$ as follows
Notice that the normalization term $\sum_{\textbf{s}\in\mathcal{S},\textbf{x}\in\mathcal{X}}\tilde{P}(\mathbf{s}|\textbf{x},\theta)$ featured in both (11) and (12), acts as a constant across all $P(Y=y|\theta)$ calculations, ensuring that $Y$ conforms to a valid probability distribution. Therefore, it only shifts the proposed UE scores which does not affect the performance of accurately predicting the correctness of the model generation. Moreover, by introducing the random variable $Y$, we not only provide a theoretical foundation for heuristic choices of the previous works but also create flexibility to define new distributions for $Y$ which may potentially improve the existing UE tools.
Using the definition of $Y$, we can also rationalize our scoring function MARS. We replace the length-normalized scoring function with MARS as in (9). We believe that MARS is a better choice to define the probability distribution of $Y$. This is because MARS considers the semantic contribution of tokens and the values of $Y$ are closely related to the semantics of the generated sentences in the context of question.
Once we do that, the new probability distribution of $P(Y=y|\theta)$ becomes the following if we consider $g$ as a one-to-one function as the work of Malinin and Gales (2021)
$P(Y=y|\theta)=\frac{\bar{P}(\mathbf{s}|\textbf{x},\theta)}{\sum_{\textbf{s}\in\mathcal{S},\textbf{x}\in\mathcal{X}}\bar{P}(\mathbf{s}|\textbf{x},\theta)}.$ (14)
If we follow Kuhn et al. (2023) and make $g$ a many-to-one function, we reach the following distribution for $P(Y=y|\theta)$:
$\begin{split}P(Y=y|\theta)&=\frac{\sum_{\textbf{s},\textbf{x}\in C_{y}}\bar{P}(\mathbf{s}|\textbf{x},\theta)}{\sum_{\textbf{s}\in\mathcal{S},\textbf{x}\in\mathcal{X}}\bar{P}(\mathbf{s}|\textbf{x},\theta)}.\end{split}$ (15)
Overall, by defining the new random variable $Y$ and the properties of meaning function $g(\cdot)$, we build a theoretical background for the heuristic design choices of previous works Malinin and Gales (2021); Kuhn et al. (2023). Moreover, this structure provides a background for further studies by either changing length-normalized scoring (as we do with MARS) or by re-defining the probability distribution of $Y$ and properties of the meaning function $g(\cdot)$.

### v3

@@ 1 Introduction
Uncertainty Estimation (UE) is a well-studied problem in classification scenarios, especially in the computer vision domain Lakshminarayanan et al. (2017); Gal and Ghahramani (2016); Shen et al. (2021). The proposed UE methods in classification tasks, which rely on the class probabilities, are not directly applicable to generative LLMs due to the auto-regressive generative structure of LLMs Malinin and Gales (2021), which implies that LLMs generate text sequentially by predicting each subsequent word based on the combined context of all preceding words. This process differs significantly from classification tasks, where the output is typically a single label or a set of labels assigned to an entire input, without the sequential and context-accumulating nature of generative LLMs.
Recent work Malinin and Gales (2021), formalizes how to adapt popular UE methods developed for classification tasks to the context of generative LLMs. They propose using length-normalized scoring to estimate the likelihood of a sequence generated by the model, and the subsequent works Kuhn et al. (2023); Lin et al. (2023); Chen and Mueller (2023) utilize that idea of length-normalized scoring.
@@
  • We explain previous works’ Malinin and Gales (2021); Kuhn et al. (2023) design choices from the classification perspective to create a grounded framework for MARS and other scoring functions.
@@ 2 Background
In this section, we will go over probability-based UE methods that our work built on. For a detailed discussion on related works, refer to Appendix A.
@@ 2.3 Length-Normalized Scoring
One of the key issues with using sequence probability $P(\mathbf{s}|\mathbf{x},\theta)$ as a proxy for $P(y|x,\theta)$ lies in its tendency to decrease as the sequence length increases. To overcome this issue, Malinin and Gales (2021) uses length-normalized scoring function instead of sequence probability.***A scoring function $K$ takes two inputs: the predicted probability $p$ of an event and its actual outcome $o$, and returns a numerical score Gneiting and Raftery (2007). Length-normalized scoring $\tilde{P}(\mathbf{s}|\mathbf{x},\theta)$ is defined as follows:
@@ 2.4 Entropy-Based UE for Generative LLMs
where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters. In Section 4, we provide an alternative explanation for Semantic Entropy and length-normalized scoring from a classification task perspective.
@@ 3.3 Importance Function Design
Our initial approach for assigning importance coefficients to tokens assumes their semantic independence even though tokens often exhibit semantic inter-dependencies. For example, in the sentence “Hamlet is written by William Shakespeare,” tokens “William” and “Shakespeare” are intrinsically linked. Treating such tokens independently ignores linguistic nuances, so we refine our methodology. Instead of masking tokens individually, we mask tokens at the phrase level (details in Appendix B.1). This approach acknowledges and preserves the inherent semantic relationships between closely related tokens, resulting in a more accurate and contextually aware assessment of token importance. In particular, a response $\textbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$ is composed of phrases $\{h_{1},h_{2},\ldots,h_{K}\}$, where each token $s_{l}$ belongs to a phrase $h_{k}$. We mask phrases one by one and find the importance coefficient of each phrase with BEM model. To translate phrase-level importance coefficients into token-level coefficients, we distribute the importance score to all tokens in the phrase equally. We summarize the enhanced algorithm in Appendix B.2. Further, in Section 5.3, we show that allocating importance score only to the most uncertain token within a phrase also yields comparable results.
@@
The necessity of performing a separate neural network pass for each phrase to determine its importance score increases the computational load of the proposed approach. Additionally, detecting phrases themselves requires another neural network pass, further increasing the computational complexity. To address these challenges, we have developed a BERT-like neural network model with 110M parameters (a significantly smaller model compared to LLMs). This model is capable of performing both tasks simultaneously for a given sequence in a single neural network pass: it identifies phrases within the generated text and their importance scores (see Figure 3). This dual-functionality substantially reduces the computational cost, making the algorithm more efficient and scalable. For detailed model architecture and performance metrics, please refer to Appendix B.
4 Understanding Generative LLM Probabilities from a Classification Perspective
In classification tasks, the class probability reflects the model’s confidence in assigning a specific class to an input. It is inherently tied to the semantics of the class. For instance, if a well-calibrated classifier gives a 75% probability to the label “cat” for a given image, it suggests a 75% likelihood that the given image is indeed a cat. This output probability is not only a numerical value; it conveys a semantic understanding of the image content as a cat. To understand previous works Kuhn et al. (2023); Malinin and Gales (2021) from a classification perspective, we propose a new random variable that is directly related to the semantics of the output.
Let $Y$ be a random variable with arbitrary dimension corresponding to the meaning of the sequences generated by an LLM parametrized with $\theta$. The values of $Y$ can be the set of all possible meanings of generated sequences and their contexts. Formally, the set is $\{g(\textbf{s},\textbf{x})\}_{\textbf{s}\in\mathcal{S},\textbf{x}\in\mathcal{X}}$, where $g(\cdot)$ is the meaning function that takes generated sentence s and context x as inputs and returns the meaning as output. A well-calibrated probability distribution of $Y$ ensures that if a model’s generation is more likely to be correct than another, the corresponding probability should be higher. Mathematically, $P(Y=g(s_{1},x_{1})|\theta)$ > $P(Y=g(s_{2},x_{2})|\theta)$ if the model’s response $s_{1}$ to question $x_{1}$ is more likely to be correct than response $s_{2}$ to question $x_{2}$. This is indeed similar to the classification example where well-calibrated class probabilities reflect the output’s actual probability of correctness Guo et al. (2017). Previous works and MARS have heuristic design choices to make $Y$ more calibrated to get a better uncertainty estimation.
Malinin and Gales (2021) considers $g(\cdot)$ as a one-to-one function which means that each unique sentence in the given context corresponds to different meanings. To achieve better calibration of $Y$ they define its distribution using length-normalized scoring of the generated sequences. This approach ensures a fair comparison between long and short sequences, thereby improving the calibration of $Y$. More formally
On the other hand, Kuhn et al. (2023) claims different sequences can have equal meaning. By considering $g(\cdot)$ as a many-to-one function, we can define the distribution of Y with their proposal as follows:
Notice that the normalization term $\sum_{\textbf{s}\in\mathcal{S},\textbf{x}\in\mathcal{X}}\tilde{P}(\mathbf{s}|\textbf{x},\theta)$ featured in both (11) and (12), acts as a constant across all $P(Y=y|\theta)$ calculations, ensuring that $Y$ conforms to a valid probability distribution. Therefore, it only shifts the proposed UE scores which does not affect the performance of accurately predicting the correctness of the model generation. Moreover, by introducing the random variable $Y$, we not only provide an alternative explanation of the previous works but also create flexibility to define new distributions for $Y$ which may potentially have better calibration and improve the existing UE tools.
Using the definition of $Y$, we can also rationalize our scoring function MARS. We replace the length-normalized scoring function with MARS as in (9). MARS considers the semantic contribution of tokens, assigning higher weights to those critical for the correctness of the answer. This emphasis on key tokens potentially makes MARS a more effective scoring function for achieving better calibration of $Y$.
@@ 5.1 Experimental Design
We use three closed-book Question-Answer (QA) datasets for evaluation: TriviaQA Joshi et al. (2017), Natural Questions Kwiatkowski et al. (2019), and WebQA Chang et al. (2022). We give further details in Appendix C.
@@
Following previous works Malinin and Gales (2021); Kuhn et al. (2023), we use Area Under the Receiver Operating Characteristic Curve (AUROC) score for our UE performance metric. AUROC quantifies a method’s ability to distinguish between two classes by plotting the true positive rate against the false positive rate for various threshold values. AUROC score is the area under this curve, ranging from 0 to 1. Higher AUROC score indicates a superior performance, while a score of 0.5 implies a random chance. In our case, ground truth is the correctnessWe use GPT-3.5-turbo for evaluating the correctness of the model, as in Lin et al. (2023); Chen and Mueller (2023). of the model response to the question and the prediction is the output of an UE method.
@@ 5.2 Main Results
We present our detailed results in Table 1. Upon closer examination of the results, it becomes apparent that the application of MARS consistently improves all baseline methods across various datasets and models. Specifically, MARS yields improvements of up to 5.8 points for Confidence, 6.24 points for Entropy, and 1.51 points for SE.
@@ 5.5 UE in Medical QA Dataset
Next, we evaluate the UE methods using a medical QA dataset. Publicly available medical QA datasets typically fall into two categories: those with multiple-choice questions Pal et al. (2022); Kotonya and Toni (2020); Jin et al. (2021) and those without clear ground truths Zhu et al. (2019, 2020). To tackle this, we create a subset from the MedMCQA multiple-choice dataset Pal et al. (2022), selecting questions that can be answered objectively without multiple choices. For this, we collaborate with medical professionals to ensure the accuracy and relevance of the selected questions, yielding a dataset of 415 samples. We use AdaptLLM’s Medicine-Chat Cheng et al. (2023), a medical-domain adapted LLaMA-2-Chat-7B modelhttps://huggingface.co/AdaptLLM/medicine-chat. To evaluate the correctness of model-generated responses, we leverage GPT-4 OpenAI (2023) and assess response validity in the medical domain.
@@ 9 Acknowledgement
This work is supported in part by a research gift from USC-Amazon Center on Secure and Trusted Machine Learning§§§https://trustedai.usc.edu . Also, we are thankful to MD. Mehmet Sahin Ulukanli and MD. Burak Yaldiz for their invaluable collaboration and expertise in the development of the Medical QA dataset.
@@ Appendix A Related Works
A.1 Discussion of the Differences with TokenSAR
There is a recent work that also considers the meaning of the words in the generation to estimate uncertainty Duan et al. (2023). The fundamental difference with our work is that Duan et al. (2023)’s method is designed as an alternative to the existing probability-based uncertainty methods, whereas, in our work, we propose a scoring function, i.e., MARS, which is compatible with all existing probability-based uncertainty estimation methods. This implies that one can in fact utilize MARS within the framework of Duan et al. (2023). In particular, in Duan et al. (2023), authors propose three schemes: TokenSAR (token-level weight assignment), SentSAR (sentence-level weight assignment), and SAR (both token and sentence-level weight assignment). SentSAR and SAR are orthogonal to MARS. SAR is the version of SentSAR where the scoring function in SentSAR is replaced with TokenSAR. In a similar fashion, MARS can be incorporated into the SentSAR approach instead of the TokenSAR.
Thus, we need to discuss our distinction from TokenSAR, which can also be considered as a scoring function. To avoid confusion and clarify our unique approach, below we discuss our distinction from TokenSAR.
  • MARS uses BERT-Matching instead of sentence similarity: In our algorithm, we utilize the BERT-Matching (BEM) model which takes the question, ground truth answer, and the generated answer as inputs and returns the probability of the generated answer being correct. To assign importance weights, we remove a set of tokens (a phrase) from the generated sentence and pass the question, generated answer (as ground truth) and token-removed generated answer to the BEM model. We set (1 - output) as the importance weight and normalize weights at the end. We further improve this process by fine-tuning a BERT-like model and increase efficiency (explained in the third bullet). TokenSAR uses sentence similarity model (cross-encoder Roberta-Large) unlike our approach. Sentence similarity model takes two input sentences to measure similarity and they concatenate the question for both inputs. However, we argue that using the BEM model achieves better performance since the goal is to find a token’s importance based on its contribution to the correctness of the generated answer. This difference becomes more visible when the answer is longer and more complex as we demonstrate in the below example. In particular, TokenSAR fails to detect words that actually answer the question so that it (almost) returns uniform importance values. On the other hand, MARS successfully finds the important words and assigns higher weights to them. Let’s consider the following example:
    Question: What is the tallest building in the world?
    Generated Answer: The Burj Khalifa in Dubai, soaring into the sky, holds the distinction of being the tallest building in the world, a marvel of modern engineering and architecture.
    To this question-answer pair, MARS returns the following importance weight assignment:
    The Burj Khalifa (0.8428) in (0.0082) Dubai (0.0083) , (0.0082) soaring (0.0084) into (0.0082) the sky (0.0082) , (0.0082) holds (0.0083) the distinction (0.0083) of (0.0082) being (0.0082) the tallest building (0.0082) in (0.0082) the world (0.0082) , (0.0082) a marvel (0.0083) of (0.0083) modern engineering and architecture (0.0088) . (0.0083)
    On the other hand, to the same pair, TokenSAR returns the following importance weight assignment:
    The (0.0225) Bur (0.0228) j (0.0318) K (0.0228) hal (0.0228) ifa (0.0319) in (0.0227) Dub (0.0253) ai (0.0232) , (0.0228) so (0.0237) aring (0.0294) into (0.0228) the (0.0228) sky (0.0235) , (0.0229) holds (0.0228) the (0.0227) distinction (0.0234) of (0.0228) being (0.0228) the (0.0229) tall (0.0235) est (0.0227) building (0.0228) in (0.0228) the (0.0228) world (0.0230) , (0.0229) a (0.0228) mar (0.0232) vel (0.0232) of (0.0230) modern (0.0540) engineering (0.0725) and (0.0336) architecture (0.0328) . (0.0232)
    In this example, although the phrase “The Burj Khalifa” is the key word answering the question, TokenSAR assigns low weights to its tokens. In fact, according to TokenSAR, tokens of the phrase “The Burj Khalifa” are as important as some of the words/phrases that appear in the question itself such as “the tallest building”. This is not ideal as TokenSAR cannot distinguish between the actual answer and filler words. However, our proposed MARS is able to actually find the important words in the answer thanks to the BEM model we employ during weight assignment.
  • MARS addresses token dependencies and process phrases instead of tokens: As we explain in Section 3, we first divide a generated answer into phrases and then assign scores to each of those phrases by using the procedure described in the first bullet point. On the other hand, Duan et al. (2023) assumes that each generated token is meaningly independent so that they remove tokens from the generation one-by-one and assign importance scores accordingly. However, as we show in Table 2, ignoring token dependencies negatively affects the performance of uncertainty methods. In this sense, our MARS provides a more careful importance score assignment (as we demonstrate in the above example).
  • MARS is computationally efficient at inference: As we mention in Section 3, we improve the computational performance of MARS by fine-tuning a BERT-like model that gives importance scores with phrases in a single forward pass (this is an improvement over the algorithm described in the first bullet). That is, our importance assignment does not depend on the number of tokens in the generated sentence. In stark contrast, Duan et al. (2023) uses cross-encoder Roberta-Large and their algorithm requires a number of tokens times forward pass for a single generated sentence. Moreover, cross-encoder Roberta-Large has approximately 355M parameters. However, we run our 110M BERT-like model only one time for the generated sentence no matter its length. That is why for a generation comprised of 10 tokens, MARS is  30x computationally more efficient than Duan et al. (2023).